# How To Segment in 3D Using 2D Models: Automated 3D Segmentation of Prostate Cancer Metastatic Lesions on PET Volumes Using Multi-Angle Maximum Intensity Projections and Diffusion Models

### v1

@@ Abstract
Prostate specific membrane antigen (PSMA) positron emission tomography/computed tomography (PET/CT) imaging provides a tremendously exciting frontier in visualization of prostate cancer (PCa) metastatic lesions. However, accurate segmentation of metastatic lesions is challenging due to low signal-to-noise ratios and variable sizes, shapes, and locations of the lesions. This study proposes a novel approach for automated segmentation of metastatic lesions in PSMA PET/CT 3D volumetric images using 2D denoising diffusion probabilistic models (DDPMs). Instead of 2D trans-axial slices or 3D volumes, the proposed approach segments the lesions on generated multi-angle maximum intensity projections (MA-MIPs) of the PSMA PET images, then obtains the final 3D segmentation masks from 3D ordered subset expectation maximization (OSEM) reconstruction of 2D MA-MIPs segmentations. Our proposed method achieved superior performance compared to state-of-the-art 3D segmentation approaches in terms of accuracy and robustness in detecting and segmenting small metastatic PCa lesions. The proposed method has significant potential as a tool for quantitative analysis of metastatic burden in PCa patients.
@@ 1 Introduction
In this work, we introduce an innovative approach for automated detection and segmentation of biochemically recurrent PCa metastatic lesions on PSMA-PET images using a 2D Diffusion-based segmentation model. This approach includes novel use of the ordered-subset expectation-maximization (OSEM) algorithm applied to 2D segmentations of multi-angle maximum intensity projections (MA-MIPs) to generate 3D segmentation of metastatic lesions in PET images, while taking advantage of the computational efficiency and performance benefits of training a 2D diffusion-based segmentation model on MA-MIPs. We show that our method outperforms its state-of-the-art 3D rivals in terms of various segmentation metrics on the target dataset.
@@ 5 Conclusion
Our work introduces a novel method for segmentation of metastatic lesions in 3D volumetric PET images, utilizing 2D diffusion models. Instead of 2D trans-axial slices or 3D volumes, our approach segments the lesions on multi-angle maximum intensity projections of PET images. Final 3D segmentation masks are then obtained through OSEM-based reconstruction of the segmented 2D MA-MIPs. Our method demonstrates superior performance compared to SOTA segmentation techniques. The proposed framework holds significant promise as a tool for segmenting small metastatic lesions in medical images.
@@ 5.0.1 Acknowledgements
This work was supported by the Canadian Institutes of Health Research (CIHR) Project Grant PJT-162216, as well as computational resources and services provided by Microsoft AI for Health.

### v3

# This manuscript has been published in Deep Generative Models. DGM4MICCAI 2024. Lecture Notes in Computer Science, vol 15224. Springer, Cham, 09 October 2024

@@ Abstract
Prostate specific membrane antigen (PSMA) positron emission tomography/computed tomography (PET/CT) imaging provides a tremendously exciting frontier in visualization of prostate cancer (PCa) metastatic lesions. However, accurate segmentation of metastatic lesions is challenging due to low signal-to-noise ratios and variable sizes, shapes, and locations of the lesions. This study proposes a novel approach for automated segmentation of metastatic lesions in PSMA PET/CT 3D volumetric images using 2D denoising diffusion probabilistic models (DDPMs). Instead of 2D trans-axial slices or 3D volumes, the proposed approach segments the lesions on generated multi-angle maximum intensity projections (MA-MIPs) of the PSMA PET images, then obtains the final 3D segmentation masks from 3D ordered subset expectation maximization (OSEM) reconstruction of 2D MA-MIPs segmentations. Our proposed method achieved superior performance compared to state-of-the-art 3D segmentation approaches in terms of accuracy and robustness in detecting and segmenting small metastatic PCa lesions. The proposed method has significant potential as a tool for quantitative analysis of metastatic burden in PCa patients. The code developed for this work is publicly available at: https://github.com/Amirhosein2c/MIP-DDPM .
@@ 1 Introduction
In this work, we introduce an innovative approach for automated detection and segmentation of biochemically recurrent PCa metastatic lesions on PSMA-PET images using a 2D Diffusion-based segmentation model. This approach includes novel use of the ordered-subset expectation-maximization (OSEM) algorithm applied to 2D segmentations of multi-angle maximum intensity projections (MA-MIPs) to generate 3D segmentation of metastatic lesions in PET images, while taking advantage of the computational efficiency and performance benefits of training a 2D diffusion-based segmentation model on MA-MIPs. We show that our method outperforms its state-of-the-art 3D rivals in terms of various segmentation metrics on the target dataset. The code developed for this work is publicly available at: https://github.com/Amirhosein2c/MIP-DDPM .
@@ 4 Results and Discussion
While our approach demonstrates promising results, it is important to consider some limitations and areas for future exploration. First, although we focused on segmentation metrics in this study, we recognize the potential value of incorporating detectability metrics, especially when dealing with small objects. This could provide additional insights into our method’s performance. The computational time of our method is an aspect that warrants further investigation. Accuracy is paramount in clinical settings, though, understanding and optimizing the time required for all the involved steps could enhance the practical applicability.
@@ 5 Conclusion
Our work introduces a novel method for segmenting small metastatic lesions on MA-MIPs using 2D DDPMs, demonstrating superior performance compared to state-of-the-art 3D segmentation methods. The proposed method represents a significant step forward in addressing the challenging clinical problem of identifying small metastatic lesions. Future directions for this research will include conducting a comprehensive ablation study to demonstrate the impact of each component in the proposed approach, exploring comparisons with other 2D and 3D diffusion-based methods, and validating on larger, multi-center datasets. By continuing to refine and expand this work, we aim to further improve the accuracy and efficiency of segmenting small objects in medical images, ultimately contributing to better patient care.
@@ 5.0.1 Acknowledgements
This work was supported by CIHR Project Grant PJT-162216, as well as computational resources provided by Microsoft AI for Health.